# Unveiling Ferroelectric-Like Behavior in Leaky Dielectrics: A Microscopic Model for Polarization Dynamics and Hysteresis Inversion

### v1

@@ II Results
Figure 1: (a) Illustration of a capapacitor with plate separation $D$ and plate area $A$ with a dielectric medium containing charge localization sites of mean size $L$. (b) Illustration of internal fluxes of nonequilibrium charges leading to polarization fluctuations. (c) Illustration of the additional presence of leaking processes as fluxes overcoming the external barriers into the sorrounding dielectric.
@@
We assume that the charge within each site redistributes either to the left or right according to the local bias, which is proportional to the total applied voltage $\eta V$, following the configuration shown in Fig. 1(a). Thus, the nonequilibrium polarization transfer described in Eq. 2 will induce a local bouncing of carriers, $\partial P/\partial t|^{(p)}$, which can be characterized in terms of carrier fluxes over left and right internal barriers. Concurrently, charges can either leak or enter the site via the external barriers, leading to a polarization fluctuation rate $\partial P/\partial t|^{(lk)}$. In a first approximation, we will assume these leaking mechanisms to be independent of the bouncing process, each with potentially distinct relaxation times, $\tau^{(p)}$ and $\tau^{(lk)}$.
@@
Here, $C=\varepsilon_{0}(1+\chi_{\infty})A/D$, $\Gamma=e\mu/D^{2}$, and $\mu$ represents the nonequilibrium carrier mobility de Paiva et al. (2022). The potential effects of drift current modulation with nonequilibrium charge fluctuations, $\delta n$, are detailed in Refs. Lopez-Richard et al. (2023a, b, 2024). Therefore, for the purposes of this study, we will assume $\Gamma\delta n<<G_{0}$, maintaining just a residual leaking conductance, $G_{0}\neq 0$, in addition to the displacement contributions. For simplicity, we have set $G_{0}=1$ as the unit of conductance in the presence of leakage.
@@
The corresponding current-voltage loops, calculated according to Eq. II, are depicted in panels 2 (e) and (f), respectively. The opening of the hysteresis response depends not only on the relative magnitude of nonequilibrium contributions ($\delta P$ in this case) compared to the stationary values, but also on the period of the cyclic voltage. Optimal conditions for a memory response, when the voltage period approaches $T\sim 2\pi\tau^{(\gamma)}$, are discussed in Refs. 26; 35; 27.
@@
For small voltage amplitudes, the saturation effects can be neglected ($\beta V^{2}\approx 0$, after Eqs. 4 and 5) and solving Eqs. 3-II in the presence of leakage, one may obtain results as those depicted in Figs 3 (a)-(d). In order to contrast the potential time-scale difference we have set $\tau^{(p)}=\tau^{(lk)}/10$ with symmetric polarization transfer ($\alpha=1$). In Fig. 3 (a), a ferroelectric-like counterclockwise loop is achieved by setting the voltage sweeping period to $T=2\pi\tau^{(p)}$, which maximizes polarization fluctuations while keeping the leakage contribution negligible. By adjusting the voltage sweep period to $T=2\pi\tau^{(lk)}$, the polarization loop direction can reverse, as shown in Fig. 3 (b). Under this condition, the voltage sweep aligns with the charge leakage dynamics, enabling the hysteresis loop to switch direction. Experimentally, clockwise reversible polarization loops have been observed in Ref. 36, and similar effects have been investigated in Ref. 37. In addition to frequency tuning, the balance between polarization transfer and charge leakage is significantly influenced by temperature, governed by the parameters $\lambda^{(p)}$ and $\lambda^{(lk)}$ as defined after Eq. 5, as confirmed in Ref. 36. The impact of temperature tuning on these effects will be further addressed in the discussion.
@@
Note that despite the reversal of the polarization loop, the current loops have consistently maintained a clockwise direction throughout both Fig. 2 and Fig. 3. This behavior arises from the relatively strong contribution of the capacitive term in Eq. II, $CdV/dt$. For a negligible capacitive contribution (with $C\rightarrow 0$), a reversal of the current loops is also expected. In Fig. 4, we compare two scenarios with parameters minimizing the capacitive term: the left column shows a larger value of $\lambda^{(p)}$, where no loop reversal is observed within the analyzed period range, while for a smaller $\lambda^{(p)}$, a counterclockwise current loop emerges as the period increases. This effect has been experimentally observed in Ref. 37 and attributed to the detrapping of electrons from interface traps, in agreement with the theoretical mechanisms described in our model.
@@
Under a sinusoidal voltage input, $V=V_{0}\cos(\omega t)$, and by expanding Eqs. 3-5 in powers of $\eta eV_{0}/(k_{B}T_{eff})$ the solution of Eq. 7 can be derived analytically. The stable response, once all transient contributions have dissipated, can be expressed as a combination of harmonics Lopez-Richard et al. (2023b)
@@
In this representation, the impedance per mode can be expressed as Lopez-Richard et al. (2023b)

### v2

@@ II Results
Figure 1: (a) Illustration of a capapacitor with plate separation $D$ and plate area $A$ with a dielectric medium containing charge localization sites of mean size $L$. (b) Barrier profile showing the carrier fluxes overcoming $E^{(p)}$ barriers that lead to the charge bouncing within the site and the carrier fluxes overcoming $E^{(lk)}$ barriers that lead to the charge leak. (c) Illustration of internal fluxes of nonequilibrium charges leading to polarization fluctuations with the presence of leaking processes as fluxes overcoming the external barriers into the sorrounding dielectric.
@@
We assume that the charge within each site redistributes either to the left or right according to the local bias, which is proportional to the total applied voltage $\eta V$, following the configuration shown in Fig. 1(b). Thus, the nonequilibrium polarization transfer described in Eq. 2 will induce a local bouncing of carriers, $\partial P/\partial t|^{(p)}$, which can be characterized in terms of carrier fluxes over left and right internal barriers. Concurrently, charges can either leak or enter the site via the external barriers, leading to a polarization fluctuation rate $\partial P/\partial t|^{(lk)}$. In a first approximation, we will assume these leaking mechanisms to be independent of the bouncing process, each with potentially distinct relaxation times, $\tau^{(p)}$ and $\tau^{(lk)}$.
@@
Here, $C=\varepsilon_{0}(1+\chi_{\infty})A/D$, $\Gamma=e\mu/D^{2}$, and $\mu$ represents the nonequilibrium carrier mobility de Paiva et al. (2022). The potential effects of drift current modulation with nonequilibrium charge fluctuations, $\delta n$, are detailed in Refs. Lopez-Richard et al. (2023, 2024a, 2024b, 2024c). Therefore, for the purposes of this study, we will assume $\Gamma\delta n<<G_{0}$, maintaining just a residual leaking conductance, $G_{0}\neq 0$, in addition to the displacement contributions. For simplicity, we have set $G_{0}=1$ as the unit of conductance in the presence of leakage.
@@
The corresponding current-voltage loops, calculated according to Eq. II, are depicted in panels 2 (e) and (f), respectively. The opening of the hysteresis response depends not only on the relative magnitude of nonequilibrium contributions ($\delta P$ in this case) compared to the stationary values, but also on the period of the cyclic voltage. Optimal conditions for a memory response, when the voltage period approaches $T\sim 2\pi\tau^{(\gamma)}$, are discussed in Refs. 26; 36; 27.
@@
For small voltage amplitudes, the saturation effects can be neglected ($\beta V^{2}\approx 0$, after Eqs. 4 and 5) and solving Eqs. 3-II in the presence of leakage, one may obtain results as those depicted in Figs 3 (a)-(d). In order to contrast the potential time-scale difference we have set $\tau^{(p)}=\tau^{(lk)}/10$ with symmetric polarization transfer ($\alpha=1$). In Fig. 3 (a), a ferroelectric-like counterclockwise loop is achieved by setting the voltage sweeping period to $T=2\pi\tau^{(p)}$, which maximizes polarization fluctuations while keeping the leakage contribution negligible. By adjusting the voltage sweep period to $T=2\pi\tau^{(lk)}$, the polarization loop direction can reverse, as shown in Fig. 3 (b). Under this condition, the voltage sweep aligns with the charge leakage dynamics, enabling the hysteresis loop to switch direction. Experimentally, clockwise reversible polarization loops have been observed in Ref. 37, and similar effects have been investigated in Ref. 38. The term proteresis has been coined Girard and Boissel (1989) to describe these kind of loops characterized by an apparent anticipation of the response Wei et al. (2008); Syed et al. (2020); Khan et al. (2021), in contrast to hysteresis (coming from the Greek meaning something that comes later Girard and Boissel (1989)), which would represent delayed dynamics. While this terminology aims to capture the opposite nature of such loops, it can lead to the misconception that the apparent anticipation suggests a lack of causality. Our model demonstrates that the direction of the loop is not related to the perturbation timing but it is fully attributable to the nature of the transfer function in Eq. 3, whether it involves carrier trapping or release, both of which are inherently delayed processes governed by specific relaxation times according to Eq. 7. For this reason, we advocate that using the term inverted hysteresis remains accurate and appropriate, contrary to suggestions made elsewhere.
In addition to frequency tuning, the balance between polarization transfer and charge leakage is significantly influenced by temperature, governed by the parameters $\lambda^{(p)}$ and $\lambda^{(lk)}$ as defined after Eq. 5, as confirmed in Ref. 37. The impact of temperature tuning on these effects will be further addressed in the discussion.
@@
Note that despite the reversal of the polarization loop, the current loops have consistently maintained a clockwise direction throughout both Fig. 2 and Fig. 3. This behavior arises from the relatively strong contribution of the capacitive term in Eq. II, $CdV/dt$. For a negligible capacitive contribution (with $C\rightarrow 0$), a reversal of the current loops is also expected. In Fig. 4, we compare two scenarios with parameters minimizing the capacitive term: the left column shows a larger value of $\lambda^{(p)}$, where no loop reversal is observed within the analyzed period range, while for a smaller $\lambda^{(p)}$, a counterclockwise current loop emerges as the period increases. This effect has been experimentally observed in Ref. 38 and attributed to the detrapping of electrons from interface traps, in agreement with the theoretical mechanisms described in our model.
@@
Under a sinusoidal voltage input, $V=V_{0}\cos(\omega t)$, and by expanding Eqs. 3-5 in powers of $\eta eV_{0}/(k_{B}T_{eff})$ the solution of Eq. 7 can be derived analytically. The stable response, once all transient contributions have dissipated, can be expressed as a combination of harmonics Lopez-Richard et al. (2024a)
@@
In this representation, the impedance per mode can be expressed as Lopez-Richard et al. (2024b, a)